# On the Trade-off between Flatness and Optimization in Distributed Learning

### v1

@@ Abstract
This paper proposes a theoretical framework to evaluate and compare the performance of gradient-descent algorithms for distributed learning in relation to their behavior around local minima in nonconvex environments. Previous works have noticed that convergence toward flat local minima tend to enhance the generalization ability of learning algorithms. This work discovers two interesting results. First, it shows that decentralized learning strategies are able to escape faster away from local minimizers and favor convergence toward flatter minima relative to the centralized solution in the large-batch training regime. Second, and importantly, the ultimate classification accuracy is not solely dependent on the flatness of the local minimizer but also on how well a learning algorithm can approach that minimum. In other words, the classification accuracy is a function of both flatness and optimization performance. The paper examines the interplay between the two measures of flatness and optimization error closely. One important conclusion is that decentralized strategies of the diffusion type deliver enhanced classification accuracy because it strikes a more favorable balance between flatness and optimization performance.
@@ I Introduction and related works
As modern society continues to evolve into a more interconnected world, and large-scale applications become increasingly prevalent, the interest in distributed learning has grown significantly [1, 2, 3]. In the distributed scenario, a group of agents, each with its own objective function, collaborate to optimize a global objective. To achieve the goal, two popular gradient descent based methodologies are commonly employed in algorithm design [4, 5, 6, 7]. The first one, known as the centralized method, requires all agents to send their data to a central server that manages all computations. In the second approach, which is called decentralized, agents are connected by a graph topology, and they process data locally and exchange information with their neighbors. Furthermore, the terms consensus and diffusion refer to two prominent decentralized methods in the optimization and machine learning communities [8, 9, 10, 1, 11, 12, 13, 14]. It is widely known that decentralized implementations offer enhanced data privacy, resilience to failure and reduced computation burden compared to the centralized approach [6, 9, 15, 16, 17, 18, 19].
@@
As for the fundamental properties of the three strategies including centralized, consensus and diffusion, the works by [6, 20, 11, 21, 15, 22, 23] provide convergence guarantees for convex and non-convex environments. The results generally suggest that decentralized operations could, at best, only match the optimization performance of the centralized approach [6, 24, 25], or potentially degrade the generalization ability of models [26, 10, 27]. Nevertheless, the work [28] empirically demonstrated that the network disagreement, i.e., the distance among agents, in the middle of decentralized training can enhance generalization over centralized training. Afterwards, the work [9] provably verified how the consensus strategy enhances the generalization performance of models over centralized solutions in the large-batch setting from the perspective of sharpness-aware minimization (SAM), which is primarily designed to reduce the sharpness of machine learning models [29]. Specifically, the results in [9] showed that the network disagreement among agents, which changes over time, during training with the consensus method implicitly introduces an extra regularization term related to SAM to the original risk function, which helps drive the iterates to flatter models that are known to generalize better. However, as Figure 1 shows, flatter models obtained by consensus do not always guarantee better test accuracy in the context of distributed learning. This is because the final test accuracy depends on both generalization and optimization performance. Although decentralized training tends to favor flatter minima that generalize better than those found by the centralized method, it could also implicitly degrade the optimization performance in some cases. This observation motivates us to examine the behavior of centralized and decentralized strategies more closely and to compare more directly both their generalization and optimization abilities. As a result, this work discovers two interesting results. First, it shows that decentralized learning strategies are able to escape faster away from local minimizers and favor convergence toward flatter minima relative to the centralized solution. Second, and importantly, the ultimate classification accuracy is not solely dependent on the flatness of the local minimizer but also on how well a learning algorithm can approach that minimum. In other words, the classification accuracy is a function of both flatness and optimization performance. The paper examines the interplay between the two measures of flatness and optimization error closely. One important conclusion is that decentralized strategies of the diffusion type deliver enhanced classification accuracy because they strike a more favorable balance between flatness and optimization performance.
@@ I-A Related works
Understanding the optimization behavior of mini-batch gradient descent (GD) algorithms and their influence on the generalization of models has emerged as a prominent topic of interest in recent years. Large-batch training is increasingly important in distributed learning for its potential to enhance the training speed and scalability of modern neural networks. However, it has already been observed in the literature that small-batch GD tends to flatter minima than the large-batch version [30], and that flat minima usually generalize better than sharp ones [31, 32, 33, 34, 35, 36]. Intuitively, the loss values around flat minima change slowly when the model parameters are adjusted, thereby reducing the disparity between the training and test data [30]. The works by [37, 38, 39, 40, 34, 41] utilized the stochastic differential equation (SDE) approach as a fundamental tool to analyze the regularization effects of mini-batch GD. This type of analysis inherently introduces an extra layer of approximation error due to substituting the discrete-time update by a different equation. Moreover, to enable the analysis with SDE, it is necessary to make some extra assumptions on the gradient noise, such as being Gaussian distributed [34, 39], parameter-independent [37] or heavy-tailed distributed [42]. To overcome these limitations, another line of works [43, 44, 33] focused on the dynamical stability of mini-batch GD with discrete-time analysis. Specifically, they determined the conditions on the Hessian matrices to ensure that the distance to a local minimum does not increase, thereby stabilizing the iterates in the vicinity of a local minimum.
@@ I-B Contribution
Contrary to the aforementioned works focusing on the single-agent case, we will study in this work the optimization bias of distributed algorithms in the multi-agent setting, which allows us to compare the performance of various distributed strategies. Inspired by [34], we start from investigating the escaping efficiency of algorithms from local minima, and then relate it to the trade-off between flatness and optimization. Our contributions are listed as follows:
@@
(1) We propose a general framework to examine the escaping efficiency of distributed algorithms from local minima by staying in the discrete-time domain. Since the dependence of the Hessian matrix on the immediate iterate in the original recursion makes the optimization analysis intractable, it is necessary to resort to a short-term model where the original Hessian is replaced by one evaluated at some local minima. We rigorously verify that the approximation error between the short-term and true models is negligible, ensuring that the short-term model represents the true model accurately enough. Afterwards, we obtain closed-form expressions of the short-term excess risk which quantifies the escaping efficiency. Note that we follow a different discrete-time analysis from the dynamical stability approach used in [43, 44, 33] and which focused on studying the stability of algorithms. Similar to [34], our emphasis is on quantifying the extent to which algorithms can escape from local minima.
@@
(2) According to the results obtained from (1), we compare the escaping efficiency of the centralized, consensus and diffusion methods. We show that decentralized approaches, i.e., consensus and diffusion, gain additional efficacy from the network heterogeneity and graph structure, enabling them to escape faster from local minimum than the centralized strategy. We further verify that consensus outperforms diffusion in terms of escaping ability from local minima. We also show that higher escaping efficiency encourages algorithms to favor flatter minimum by relating escaping efficiency to the flatness metric.
@@
(3) If the additional power generated by the network heterogeneity and graph structure are not sufficiently strong to allow decentralized methods to successfully leave the current basin, then both decentralized and centralized methods will be stuck within the basin of the current local minimum. A similar rule also holds for the comparison between consensus and diffusion. This motivates us to pursue next a long-term analysis, which corresponds to the optimization performance. In this context, we verify that the extra power boosting the escaping efficiency could inversely deteriorate the optimization performance. This reveals the inherent trade-off between flatness and optimization, with diffusion showing superior optimization performance over consensus in terms of smaller estimation error.
(4) We finally illustrate the performance of centralized, consensus and diffusion training strategies on real data. In addition to the flatness and optimization performance of the three distributed methods, we show that diffusion achieves a favorable balance between flatness and optimization, thereby exhibiting superior test accuracy than the other two methods (i.e., centralized and consensus).
@@ II-A Empirical risk minimization
Since we aim to understand the optimization behavior of algorithms around local minima of nonconvex risk functions, we need to distinguish between the local minima of $J_{k}(w)$ and $J(w)$. Thus, we let $w^{\star}$ denote a local minimizer for the aggregate risk $J(w)$, and let $w_{k}^{\star}$ denote a local minimizer for the individual risk $J_{k}(w)$. If all agents in the network share the same risk function, i.e., $J_{k}(w)=J(w)$, then $w^{\star}=w_{k}^{\star}$, and we refer to the network as being homogeneous. In this case, all local minimizers of $J(w)$ are also local minima of $J_{k}(w)$. Otherwise, we refer to the network as being heterogeneous, where the $w_{k}^{\star}$ of all agents can be distinct among themselves and also in relation to $w^{\star}$. In this paper, we focus on this latter more general case.
II-B Gradient-descent algorithms
@@
$\boldsymbol{w}_{n}=\boldsymbol{w}_{n-1}-\mu\times\frac{1}{KB}\sum\limits_{k=1}^{K}\sum\limits_{b=1}^{B}Q_{k}(\boldsymbol{w}_{n-1};\boldsymbol{x}_{k,n}^{b})$ (3)
Here, $B$ is the batch size and $\mu$ is a small positive step-size parameter. Observe that we are using boldface letters to refer to the data samples and iterates in (3); it is our convention in this paper to use the boldface notation to refer to random quantities.
@@ II-C Escaping efficiency from local minima
(Basin of attraction [38].) For a given local minimizer $w^{\star}$, its basin of attraction $\Omega(w^{\star})$ (or the valley of $w^{\star}$) is defined as the set of all points starting from which $w_{k,n}$ or $w_{n}\to w^{\star}$ as $n\to\infty$ if the step size $\mu$ is sufficiently small and there is no noise in the gradient-descent algorithms.
@@
We illustrate the basin of attraction related to $w^{\star}$ in Figure 2, where $\Omega(w^{\star})=(w_{1},w_{2})$ is a bounded open set. Also, we use $\partial\Omega(w^{\star})$ to denote the boundary of $\Omega(w^{\star})$, according to which $\partial\Omega(w^{\star})=\{w_{1},w_{2}\}$ in Figure 2. If ${w}_{k,n}$ or $w_{n}\notin\Omega(w^{\star})$, then we say the algorithm escapes from $w^{\star}$ exactly.
Figure 2: An illustration of the valley around local minimizer $w^{\star}$.
@@
The larger this value is, the higher the escaping efficiency will be away from $w^{\star}$. If the algorithm ultimately converges to $w^{\star}$, then the larger $\mathrm{ER}_{n}$ is, the worse the optimization performance will be.
@@
$\mathrm{ER}_{n}$ assesses the average excess-risk value across the network, and it quantifies the deviation between the current model at iteration $n$ and the local minimum. For a fixed $n$, a larger value of $\mathrm{ER_{n}}$ indicates an expected faster escape from the local minimum (since the iterate will be further away from it). To justify whether an algorithm escapes from a local basin or not, earlier studies [46, 47, 48] have established various criteria based on the analysis methods they use. They nevertheless generally adhered to the principle of the minimum effort to reach the boundary of the local basin. In a similar vein, we introduce the risk barrier, which is defined as the infimum of the risk gap between the boundary points and the local minimum:
@@
For example, the risk barrier of $w^{\star}$ in Figure 2 is $J(w_{2})-J(w^{\star})$. When $\mathrm{ER}_{n}$ exceeds the corresponding risk barrier defined in (8), namely,
$\mathrm{ER}_{n}\geq h$ (9)
then we say that the network escapes from $w^{\star}$ on average. If, on the other hand, $\boldsymbol{w}_{k,n}$ gets stuck around (or converges to) $w^{\star}$, then $\mathrm{ER}_{n}$ in that case would serve as a measure of the resulting optimization performance[6, 7].
@@ II-D Flatness metrics
Next, we formally characterize the notion of flat minima. Consider a local minimum $w^{\star}$ and a small drift $\Delta w$ around it. The change in the risk value can be approximated by
$J(w^{\star}+\Delta w)-J(w^{\star})\approx\|\Delta w\|^{2}_{H^{\star}}$ (10)
where $H^{\star}$ refers to the(positive definite) Hessian matrix of $J(w)$ at $w^{\star}$. If the above change in the risk value is small then we refer to $w^{\star}$ as a flat minimum. Otherwise, if the change in the risk value is large, then we refer to $w^{\star}$ as a sharp minimum. Motivated by (10), various metrics related to the Hessian matrix are applied in the literature to measure the flatness of local minima, e.g., the spectral norm $\|H^{\star}\|_{2}$[33, 43], the Frobenius norm $\|H^{\star}\|_{F}$[33, 44] and the trace $\mathrm{Tr}(H^{\star})$[49, 33, 50]. In this paper, we use the trace of the Hessian as a flatness measure, but the other two metrics may also be applied due to the norm equivalence, namely:

### v2

@@ Abstract
This paper proposes a theoretical framework to evaluate and compare the performance of stochastic gradient algorithms for distributed learning in relation to their behavior around local minima in nonconvex environments. Previous works have noticed that convergence toward flat local minima tend to enhance the generalization ability of learning algorithms. This work discovers three interesting results. First, it shows that decentralized learning strategies are able to escape faster away from local minima and favor convergence toward flatter minima relative to the centralized solution. Second, in decentralized methods, the consensus strategy has a worse excess-risk performance than diffusion, giving it a better chance of escaping from local minima and favoring flatter minima. Third, and importantly, the ultimate classification accuracy is not solely dependent on the flatness of the local minimum but also on how well a learning algorithm can approach that minimum. In other words, the classification accuracy is a function of both flatness and optimization performance. In this regard, since diffusion has a lower excess-risk than consensus, when both algorithms are trained starting from random initial points, diffusion enhances the classification accuracy. The paper examines the interplay between the two measures of flatness and optimization error closely. One important conclusion is that decentralized strategies deliver in general enhanced classification accuracy because they strike a more favorable balance between flatness and optimization performance compared to the centralized solution.
@@ I Introduction and related works
As modern society continues to evolve into a more interconnected world, and large-scale applications become increasingly prevalent, the interest in distributed learning has grown significantly [1, 2, 3]. In the distributed scenario, a group of agents, each with its own objective function, collaborate to optimize a global objective. To achieve the goal, two popular gradient descent based methodologies are commonly employed in algorithm design [4, 5, 6, 7]. The first one, known as the centralized method, requires all agents to send their data to a central server that manages all computations. In the second approach, which is called decentralized, agents are connected by a graph topology, and they process data locally and exchange information with their neighbors. It is widely known that decentralized implementations offer enhanced data privacy, resilience to failure and reduced computation burden compared to the centralized approach [6, 8, 9, 10, 11, 12, 13]. Consensus and diffusion strategies are two prominent decentralized methods in the optimization and machine learning communities [14, 8, 15, 1, 16, 17, 18, 19].
@@
Extensive works exist in the literature on the fundamental properties of the three strategies including centralized, consensus, and diffusion. For example, [6, 7, 16, 20, 9, 21, 22] provided convergence guarantees for convex and non-convex environments. The results generally suggest that decentralized operations could, at best, only match the optimization performance of the centralized approach [6, 23, 24], or potentially degrade the generalization ability of models [25, 15, 26]. Nevertheless, the work [27] empirically demonstrated that the network disagreement, i.e., the distance among agents, in the middle of decentralized training can enhance generalization over centralized training. Afterwards, the work [8] verified that the consensus strategy enhances the generalization performance of models over centralized solutions in the large-batch setting from the perspective of sharpness-aware minimization (SAM), which is primarily designed to reduce the sharpness of machine learning models [28]. Specifically, the results in [8] showed that the network disagreement among agents, which changes over time during training with the consensus method, implicitly introduces an extra regularization term related to SAM to the original risk function, which helps drive the iterates to flatter models that are known to generalize better. However, as Figure 1 shows, flatter models obtained by consensus do not always guarantee better test accuracy in the context of distributed learning. This is because the final test accuracy depends on both generalization and optimization performance. Although decentralized training tends to favor flatter minima that generalize better than those found by the centralized method, it could also implicitly degrade the optimization performance in some cases. This observation motivates us to examine the behavior of centralized and decentralized strategies more closely and to compare more directly both their generalization and optimization abilities.
As a result, this work discovers three interesting results. First, it shows that decentralized learning strategies are able to escape faster away from local minima and favor convergence toward flatter minima relative to the centralized solution. Second, in decentralized methods, the consensus strategy has a worse excess-risk performance than diffusion, giving it a better chance of escaping from local minima and favoring flatter minima. Third, and importantly, the ultimate classification accuracy is not solely dependent on the flatness of the local minimum but also on how well a learning algorithm can approach that minimum. In other words, the classification accuracy is a function of both flatness and optimization performance. In this regard, diffusion has a better optimization performance than consensus. Therefore, when both algorithms are trained starting from random initial points, diffusion leads to better accuracy. The paper examines the interplay between the two measures of flatness and optimization error closely. One important conclusion is that decentralized strategies deliver in general enhanced classification accuracy because they strike a more favorable balance between flatness and optimization performance compared to the centralized solution. Moreover, consensus and diffusion exhibit different trade-off abilities between optimization and flatness across different training scenarios. Due to space limitations, complete proofs for all lemmas and theorems are provided in the supplementary files.
@@ I-A Related works
Understanding the optimization behavior of mini-batch stochastic gradient algorithms (SGA) and their influence on the generalization of models has emerged as a prominent topic of interest in recent years. Large-batch training is increasingly important in distributed learning for its potential to enhance the training speed and scalability of modern neural networks. However, it has been observed in the literature that small-batch SGA tends to flatter minima than the large-batch version [29], and that flat minima usually generalize better than sharp ones [30, 31, 32, 33, 34, 35]. Intuitively, the loss values around flat minima change slowly when the model parameters are adjusted, thereby reducing the disparity between the training and test data [29]. The works by [36, 37, 38, 39, 33, 40] utilized the stochastic differential equation (SDE) approach as a fundamental tool to analyze the regularization effects of mini-batch SGA. This type of analysis inherently introduces an extra layer of approximation error due to substituting the discrete-time update by a differential equation. Moreover, to enable the analysis with SDE, it is necessary to introduce some extra assumptions on the gradient noise, such as being Gaussian distributed [33, 38], parameter-independent [36] or heavy-tailed distributed [41]. To overcome these limitations, another line of work [42, 43, 32] focused on the dynamical stability of mini-batch SGA with discrete-time analysis. Specifically, they determined conditions on the Hessian matrices to ensure that the distance to a local minimum does not increase, thereby stabilizing the iterates in the vicinity of a local minimum.
@@ I-B Contribution
Contrary to the aforementioned works focusing on the single-agent case, we will study in this work the optimization bias of distributed algorithms in the multi-agent setting, which allows us to compare the performance of various distributed strategies. Inspired by [33], we start from investigating the escaping efficiency of algorithms in the vicinity of a local minimum, and then relate it to the trade-off between flatness and optimization. Our contributions are listed as follows:
@@
(1) We propose a general framework to examine the escaping efficiency of distributed algorithms from a local minimum while remaining in the discrete-time domain. Since the dependence of the Hessian matrix on the immediate iterate in the original recursion makes the optimization analysis intractable, it is necessary to resort to a short-term model where the original Hessian is replaced by one evaluated at a local minimum. We rigorously verify that the approximation error between the short-term and true models is negligible, ensuring that the short-term model represents the true model accurately enough. Afterwards, we obtain closed-form expressions of the short-term excess risk which quantifies the escaping efficiency. Note that we follow a different discrete-time analysis from the dynamical stability approach used in [42, 43, 32] that focused on studying the stability of algorithms. Our emphasis is on quantifying the extent to which algorithms can escape local minima.
@@
(2) We compare the escaping efficiency of the centralized, consensu, and diffusion methods. Our analysis reveals that decentralized approaches–namely, consensus and diffusion–gain additional efficacy from the network heterogeneity and graph structure, making them more efficient in escaping a local minimum than the centralized strategy. Also, we show that the consensus strategy has a better chance of escaping from local minima than diffusion. Furthermore, we establish a connection between escaping efficiency and the flatness metric, showing that higher escaping efficiency encourages algorithms to favor flatter minima. This suggests that decentralized methods are more likely to converge to flatter minima compared to their centralized counterpart.
@@
(3) If the additional power generated by the network heterogeneity and graph structure are not sufficiently strong to allow decentralized methods to successfully leave the current basin, then both decentralized and centralized methods will be stuck within the basin of the current local minimum. A similar rule also holds for the comparison between consensus and diffusion. This motivates us to pursue next a long-term analysis, which corresponds to the optimization performance, i.e., how close algorithms approach the local minimum. In this context, we verify that the extra power boosting the escaping efficiency could inversely deteriorate the optimization performance. This reveals an inherent trade-off between flatness and optimization.
(4) We finally illustrate the performance of centralized, consensus and diffusion training strategies on real data. Basically, favoring flatter minima alone does not necessarily translate into a better classification accuracy. The accuracy is also dependent on the optimization performance of an algorithm. We show that decentralized methods in general achieve a more favorable balance between flatness and optimization compared to the centralized algorithm, thereby exhibiting better test accuracy. Furthermore, a closer comparison between consensus and diffusion reveals distinct behavior across different training scenarios. In the typical scenario of training from scratch, diffusion tends to lead to better accuracy because it provides a more favorable balance between flatness and optimization performance. However, when the decentralized methods are initialized near a local minimizer pretrained by the centralized algorithm, the “worse” optimization performance by consensus is alleviated and its classification accuracy becomes better.
@@ II-A Empirical risk minimization
Since we aim to understand the optimization behavior of algorithms around local minima of nonconvex risk functions, we need to distinguish between the local minima of $J_{k}(w)$ and $J(w)$. Thus, we let $w^{\star}$ denote a local minimizer for the aggregate risk $J(w)$ (i.e., meaning $w^{\star}$ admits a local minimum $J(w^{\star})$), and let $w_{k}^{\star}$ denote a local minimizer for the individual risk $J_{k}(w)$. If all agents in the network share the same risk function $J_{k}(w)=J(w)$, then $w^{\star}=w_{k}^{\star}$, and we refer to the network as being homogeneous. In this case, all local minimizers of $J(w)$ also admit local minima of $J_{k}(w)$. Otherwise, we refer to the network as being heterogeneous, where the $w_{k}^{\star}$ of all agents can be distinct among themselves and also in relation to $w^{\star}$. In this paper, we focus on this latter more general case.
II-B Stochastic gradient algorithms
@@
$\boldsymbol{w}_{n}=\boldsymbol{w}_{n-1}-\mu\times\frac{1}{KB}\sum\limits_{k=1}^{K}\sum\limits_{b=1}^{B}\nabla_{w}Q_{k}(\boldsymbol{w}_{n-1};\boldsymbol{x}_{k,n}^{b})$ (3)
Here, $B$ is the batch size and $\mu$ is a small positive step-size parameter (also called learning rate). Observe that we are using boldface letters to refer to the data samples and iterates in (3); it is our convention in this paper to use the boldface notation to refer to random quantities.
@@ II-C Escaping efficiency from local minima
(Basin of attraction [37].) For a given local minimizer $w^{\star}$ (or the corresponding local minimum $J(w^{\star})$, its basin of attraction $\Omega(w^{\star})$ (or the valley of $w^{\star}$) is defined as the set of all points starting from which $\boldsymbol{w}_{k,n}$ or $\boldsymbol{w}_{n}\to w^{\star}$ as $n\to\infty$ if the step size $\mu$ is sufficiently small and there is no noise in the gradient-descent algorithms.
@@
We illustrate the basin of attraction related to $w^{\star}$ in Figure 2, where $\Omega(w^{\star})=(w_{1},w_{2})$ is a bounded open set. Also, we use $\partial\Omega(w^{\star})$ to denote the boundary of $\Omega(w^{\star})$, according to which $\partial\Omega(w^{\star})=\{w_{1},w_{2}\}$ in Figure 2. If $\boldsymbol{w}_{k,n}$ or $\boldsymbol{w}_{n}\notin\Omega(w^{\star})$, then we say the algorithm escapes from the basin of $w^{\star}$ exactly.
Figure 2: An illustration of the valley around a local minimizer $w^{\star}$.
@@
We wish to highlight two aspects in relation to the definition of the escaping efficiency of algorithms:
First, $\mathrm{ER}_{n}$ quantifies the ability of the network model to escape $w^{\star}$ on average. Basically, $\mathrm{ER}_{n}$ assesses the excess-risk value across the graph on average, serving as a metric for the deviation between the excess risk at the current model at iteration $n$ and the local minimum. As Figure 2 shows, within the local basin $(w_{1},w_{2})$, a larger excess risk corresponds to a greater distance from the local minimum $J(w^{\star})$ for any single model $w$. Since this paper focuses on network models composed of $K$ local agents, we analyze the mean deviation across the network and use (7). Therefore, for a fixed $n$, a larger value for $\mathrm{ER_{n}}$ implies an expected faster escape from a local minimum (since the iterates will be further away from it).
Second, a larger value for $\mathrm{ER}_{n}$ suggests a greater likelihood to leave $\Omega(w^{\star})$. To justify whether an algorithm escapes from a local basin or not, earlier studies [46, 47, 48] have established various criteria based on the analysis methods they use. They nevertheless generally adhered to the principle of the minimum effort to reach the boundary of the local basin. In a similar vein, we introduce the risk barrier, which is defined as the infimum of the risk gap between the boundary points and the local minimum:
@@
For example, the risk barrier of $w^{\star}$ in Figure 2 is $J(w_{2})-J(w^{\star})$. For a network model, a larger $\mathrm{ER}_{n}$ reflects a higher average deviation from the local minimum, making escape more probable. Specifically, when $\mathrm{ER}_{n}$ exceeds the corresponding risk barrier defined in (8), namely, $\mathrm{ER}_{n}\geq h$, the network exits $\Omega(w^{\star})$ on average.
Therefore, a larger $\mathrm{ER_{n}}$ indicates higher escaping efficiency. If, on the other hand, the network gets stuck around (or converges to) $w^{\star}$, then $\mathrm{ER}_{n}$ in that case would serve as a measure of the resulting optimization performance[6, 7].
@@ II-D Flatness metrics
Next, we formally characterize the notion of flat minima. Consider a local minimizer $w^{\star}$ and a small drift $\Delta w$ around it. The change in the risk value can be approximated by
$J(w^{\star}+\Delta w)-J(w^{\star})\approx\frac{1}{2}\|\Delta w\|^{2}_{H^{\star}}$ (9)
where $H^{\star}$ refers to the(positive definite) Hessian matrix of $J(w)$ at $w^{\star}$. If the above change in the risk value is small then we say $w^{\star}$ admits a flat minimum. Otherwise, if the change in the risk value is large, then we say $w^{\star}$ admits a sharp minimum. Motivated by (9), various metrics related to the Hessian matrix are applied in the literature to measure the flatness of local minima, e.g., the spectral norm $\|H^{\star}\|_{2}$[32, 42], the Frobenius norm $\|H^{\star}\|_{F}$[32, 43] and the trace $\mathrm{Tr}(H^{\star})$[49, 32, 50]. In this paper, we use the trace of the Hessian as a flatness measure, but the other two metrics may also be applied due to the norm equivalence, namely: